### 21 Dec 2023 v3

# Foundation Models in Smart Agriculture: Basics, Opportunities, and Challenges

Abstract
The past decade has witnessed the rapid development and adoption of machine learning (ML) and deep learning (DL) methodologies in agricultural systems, showcased by great successes in applications such as weed control, plant disease detection, agricultural robotics, and precise livestock management. However, these conventional ML/DL models have certain limitations: They heavily rely on large, costly-to-acquire labeled datasets for training, require specialized expertise for development and maintenance, and are mostly tailored for specific tasks, thus lacking generalizability. Recently, large pre-trained models, also known as foundation models (FMs), have demonstrated remarkable successes in language, vision, and decision-making tasks across various domains. These models are trained on a vast amount of data from multiple domains and modalities. Once trained, they can accomplish versatile tasks with just minor fine-tuning and minimal task-specific labeled data. Despite their proven effectiveness and huge potential, there has been little exploration of applying FMs to agriculture artificial intelligence (AI). Therefore, this study aims to explore the potential of FMs in the field of smart agriculture. In particular, we present conceptual tools and technical background to facilitate the understanding of the problem space and uncover new research directions in this field. To this end, we first review recent FMs in the general computer science (CS) domain and categorize them into four categories: language FMs, vision FMs, multimodal FMs, and reinforcement learning FMs. Subsequently, we outline the process of developing agriculture FMs (AFMs) and discuss their potential applications in smart agriculture. We also discuss the unique challenges and risks associated with developing AFMs, including model training, validation, and deployment. Through this study, we contribute to the advancement of AI in agriculture by introducing AFMs as a promising paradigm that can significantly mitigate the reliance on extensive labeled datasets and enhance the efficiency, effectiveness, and generalization of agricultural AI systems. To facilitate further research, a well-classified and actively updated list of papers on AFMs are organized and accessible at https://github.com/JiajiaLi04/Agriculture-Foundation-Models.
keywords:
[…]
1 Introduction
Figure 1: Overview of building foundation models for agricultural applications.
Modern agricultural systems have been undergoing a revolution through the tight integration of recent advances in information and communication technologies (ICT). This has led to a prosperous field known as smart farming (also referred to as smart agriculture) (Gondchawar et al., 2016, Wolfert et al., 2017, Walter et al., 2017), showing promise in enhancing agricultural outputs, increasing farming efficiency, and improving the quality of the final product. These technologies feature multi-disciplinary advancements involving unmanned aerial/ground vehicles (UAVs/UGVs), image processing, machine learning, big data, cloud computing, and wireless sensor networks (WSNs) (Walter et al., 2017, Wolfert et al., 2017, Moysiadis et al., 2021), enabling farmers to make well-informed decisions regarding planting, tending, and harvesting to maximize productivity and profits. However, efficient extraction of relevant information from diverse data sources, especially imaging data, presents a significant challenge. Traditional data mining techniques often struggle to uncover meaningful insights from such complex data (Wolfert et al., 2017).
Deep learning (DL) (LeCun et al., 2015), on the other hand, has demonstrated remarkable capabilities in processing complex, high-dimensional data and has achieved numerous successes in a myriad of applications. DL methods excel in feature extraction, pattern classification, and learning high-quality image representations, showcasing promising performance across various agricultural domains (Kamilaris and Prenafeta-Boldú, 2018). These applications include weed control (Hasan et al., 2021, Chen et al., 2022), plant disease detection (Mohanty et al., 2016, Xu et al., 2021), plant phenotyping (Li et al., 2020, Xiang and Wang, 2023), postharvest quality assessment (Mendigoria et al., 2021, Zhou et al., 2022), and robotic fruit harvesting (Chu et al., 2021, Zhang et al., 2022a, Chu et al., 2023), among others. Despite the progress, these approaches heavily rely on supervised training, which hinges on large-scale, task-specific, and high-quality labeled datasets (Sun et al., 2017). However, collecting and annotating such datasets is extremely time-consuming, resource-intensive, and expensive, creating a high bar for resource-limited applications (Lu and Young, 2020, Xu et al., 2023). This challenge is particularly pronounced in specific applications such as plant disease detection, weed recognition, and fruit defect detection, where constraints related to biological materials, imaging conditions, and the need for precise pixel-level annotations arise (Lu and Young, 2020). Moreover, the collected datasets are often limited in their generalizability to even similar agricultural domains or applications, necessitating repetition in the data collection and model development process (Ghazi et al., 2017, Thenmozhi and Reddy, 2019). This repetitive process not only adds to the overall time and monetary costs but also hinders the generalizability, efficiency, and scalability of DL frameworks in agricultural applications.
To mitigate the aforementioned issues, many approaches have been proposed, including transfer learning (Weiss et al., 2016, Zhuang et al., 2020), few-shot learning (Wang et al., 2020, Yang et al., 2022a), label-efficient learning (Zhou, 2018, Li et al., 2023a), among others. Specifically, transfer learning has emerged as a widely adopted approach in agricultural applications and beyond (Espejo-Garcia et al., 2020, Chen et al., 2022, Abbas et al., 2021, Dang et al., 2022), which leverages DL models pre-trained on large-scale image datasets, such as ImageNet (Deng et al., 2009), Microsoft COCO (Lin et al., 2014), and PlantCLEF2022 (Goëau et al., 2022). The pre-trained models are then fine-tuned for specific tasks. For instance, Chen et al. (2022) evaluated 35 state-of-the-art DL models for weed classification using transfer learning. Few-shot learning, on the other hand, focuses on training models to quickly adopt new tasks or domains using only a limited number of labeled examples, by leveraging prior knowledge or meta-learning approaches. In contrast, label-efficient learning (Li et al., 2023a) employs weak supervision or no supervision techniques to mitigate the effects of laborious and time-consuming labeling challenges. However, these methods are often pre-trained on data from a single domain and modality, which restricts their adaptability to other domains and applications (Raffel et al., 2020). The above observations highlight the importance of incorporating diverse data sources to enhance their generalization capabilities across different domains and applications.
Large pre-trained models also referred to as foundation models (FMs) (Bommasani et al., 2021, Moor et al., 2023, Mai et al., 2023), AI models designed to produce a wide and general variety of outputs, undergo extensive training on vast and diverse datasets, enabling them to tackle numerous downstream tasks, such as speech recognition and natural language understanding (Radford et al., 2019, Brown et al., 2020, OpenAI, 2022, Ouyang et al., 2022, Touvron et al., 2023), computer vision (Rombach et al., 2022, Ramesh et al., 2022, Kirillov et al., 2023, Kang et al., 2023), and decision making (Reed et al., 2022, Team et al., 2023). FMs are generally trained on large-scale datasets from diverse domains and modalities through self-supervised learning (Liu et al., 2021), and are capable of handling a wide range of applications from various domains with minimum fine-tuning and no/little task-specific labeled data (Bommasani et al., 2021, Moor et al., 2023, Mai et al., 2023). For example, the Generative Pre-trained Transformer 3 language model (GPT-3, (Brown et al., 2020)) demonstrated a breakthrough in in-context learning, enabling the model to perform novel tasks without explicit training, by leveraging text explanations (known as “prompts”) containing just a few examples. Subsequently, ChatGPT (OpenAI, 2022), a derivative of GPT-3, was developed and publicized by OpenAI for chat-based interactions and has brought about a transformative evolution in natural language processing, enabling unprecedentedly engaging conversational experiences. In the computer vision domain, FMs such as the segment anything model (SAM, (Kirillov et al., 2023)), trained on over 1 billion masks on 11M licensed and privacy-respecting images, can solve a range of downstream segmentation problems with new image distributions and tasks with zero-shot generalization. Deepmind Ada (Team et al., 2023), equipped with a customized Transformer architecture (Vaswani et al., 2017) using distillation with a teacher-student approach (Schmitt et al., 2018), brought FMs to reinforcement learning (RL) that generalizes to new tasks in a few-shot setting. Despite the aforementioned advances, FMs’ adoption in agriculture artificial intelligence (AI) has received scarce attention.
As such, this study aims to explore the potential of developing and applying FMs for agricultural applications. More specifically, we first review recent FMs (Section 2) in the general computer science (CS) domain and categorize them into four categories: language FMs, vision FMs, multimodal FMs, and reinforcement learning FMs. Subsequently, we outline the process of developing AFMs (Figure 1) and discuss their potential applications (Section 3) in smart agriculture. In addition, this study also identifies and discusses the unique challenges and potential pitfalls associated with developing AFMs (Section 4), including model training, validation, and deployment. This comprehensive examination of AFMs can be served as a valuable resource for newcomers to the field, as well as experienced researchers seeking to innovate within the agricultural space.
2 Foundation Models (FMs)
In this section, we first cover both the strengths and weaknesses of FMs, giving readers a balanced view of the topic. Then, we explore the different kinds of FMs, their distinct characteristics, and their applicability to various domains or tasks.
2.1 Pros and Cons of FMs
FMs represent a significant paradigm shift in the field of artificial intelligence (AI). In contrast to traditional AI models that are trained on domain-specific data for specialized tasks, FMs are pre-trained on a vast scale of data, spanning multiple domains and modalities (Bommasani et al., 2021, Moor et al., 2023, Mai et al., 2023). The training process allows these models to learn a broad base of knowledge and skills, much like the foundation of a building that supports all the structures above it. Once trained, these FMs can accomplish versatile tasks with just a little fine-tuning. This is analogous to the way humans transfer knowledge and skills from one area to another (Moor et al., 2023). They have led to breakthroughs in a variety of fields, from natural language processing to computer vision and beyond. We can summarize the pros of FMs as follows:
  1. 1.
    Pre-trained Knowledge: By training on vast and diverse datasets, FMs possess a form of “general intelligence” that encompasses knowledge of the world, language, vision, and their specific training domains.
  2. 2.
    Fine-tuning Flexibility: FMs demonstrate superior performance to be fine-tuned for particular tasks or datasets, saving the computational and temporal investments required to train extensive models from scratch.
  3. 3.
    Data Efficiency: FMs harness their foundational knowledge, exhibiting remarkable performance even in the face of limited task-specific data, which is effective for scenarios with data scarcity issues.
Given the distinct advantages of FMs over traditional AI models, they exhibit substantial promise for agricultural applications. However, despite their progress, FMs are not without their unique set of challenges:
  1. 1.
    Computational Cost: The foundational training of these models demands considerable computational resources, often beyond the reach of individuals or firms with constrained budgets.
  2. 2.
    Generalization v.s. Specialization: Although FMs offer commendable generalization capabilities across diverse tasks, there exist scenarios where task-specific models deliver superior performance.
  3. 3.
    Deployment Difficulties: The considerable size of many FMs poses challenges for deployment on edge devices or other systems with restricted computational and storage capabilities.
  4. 4.
    Transparency and Interpretability: The complexity and scale of FMs can hinder clear understanding and interpretation, making them opaque in certain applications.
  5. 5.
    Ethical Concerns: Given their expansive training datasets, there’s potential for these models to inadvertently perpetuate and magnify existing societal biases.
For an in-depth discussion on deploying FMs in agricultural contexts, please refer to Section  4.
2.2 Categories of FMs
In this subsection, we present comprehensive reviews of FMs within the computer science (CS) domain, which were categorized into four distinct types (Figure 2) based on the primary data modality and learning paradigm they address: language FMs (LFMs), vision FMs (VFMs), multimodal FMs (MFMs), and reinforcement learning (RLFMs).
Figure 2: Timeline of foundation models (FMs).
2.2.1 Language foundation models (LFMs)
Over the past few years, the field of computational natural language has seen a complete transformation (Devlin et al., 2018, Brown et al., 2020) due to the rise of large-scale language modeling (LLMs). Essentially, language modeling (Nadkarni et al., 2011) involves predicting the subsequent element in a given sequence of tokens, relying on a self-supervised objective that requires no manual labeling beyond a raw text corpus.
The Bidirectional Encoder Representations from Trans-
formers (BERT, (Devlin et al., 2018)) is one of the most prominent models, with two versions: $\text{BERT}_{\text{BASE}}$ and
$\text{BERT}_{\text{LARGE}}$, containing 110M and 340M parameters respectively. By simply incorporating an additional output layer for fine-tuning, the pre-trained BERT model can be adapted to generate cutting-edge models for a variety of tasks, such as question answering and language inference. Remarkably, this can be achieved without requiring substantial modifications to the task-specific architecture, even surpassing the performance of many dedicated task-specific architectures. Subsequently, Google proposed the Pathways Language Model (PaLM, (Wei et al., 2022)) and LLAMA (Touvron et al., 2023), further advancing the field of natural language processing. PaLM, with 540B parameters, was designed with a dense decoder-only Transformer model trained with the Pathways system (Barham et al., 2022) to enable efficient training. PaLM surpassed few-shot performance of prior large models, such as GLaM (Du et al., 2022), GPT-3 (Brown et al., 2020), and LLAMA (Touvron et al., 2023), on numerous very difficult tasks, such as nature language inference (Bowman et al., 2015), in-context reading comprehension (Jenkins et al., 2003), and question answering (Kwiatkowski et al., 2019).
Participating in the development of large language models, OpenAI has proposed noteworthy contributions such as Generative Pretrained Transformer 2 (GPT-2, (Radford et al., 2019)), 3 (GPT-3, (Brown et al., 2020)), and 4 (GPT-4, (OpenAI, 2023)), having grown progressively more powerful, with 1.5 billion, 175 billion, and a staggering number of parameters, respectively, and setting remarkable benchmarks in language processing. ChatGPT (OpenAI, 2022), a specialized version of these models, has been extensively fine-tuned for human-like conversation. By adjusting the final output layer, the pre-trained GPT models can be tailored to deliver exceptional performance across a wide range of tasks, such as text generation, translation, and dialogue systems, outperforming many task-specific models without requiring extensive modifications to their foundational architecture. Recently, OpenAI showed that InstructGPT (Ouyang et al., 2022) with only 1.3B parameters are much better at following users’ instructions than GPT-3, by using an existing technique called reinforcement learning from human feedback (RLHF, (Christiano et al., 2017)).
With the continuous growth in data availability and computing power, language modeling has become a trustworthy method for crafting ever more powerful models.
2.2.2 Vision foundation models (VFMs)
Computer vision, a field dedicated to understanding and analyzing visual data, has traditionally relied on task-specific models tailored to specific objectives. These models, such as ResNet (He et al., 2016) for image classification, Faster RCNN (Ren et al., 2015) for object detection, Detectron (Girshick et al., 2018) for instance segmentation, and DCGAN (Radford et al., 2015) for image generation, have made significant contributions to advancing visual perception tasks. However, it is important to note that these task-specific models are designed to excel in their respective applications and may not be directly applicable to other computer vision tasks (Bommasani et al., 2021).
With the emergence of vision FMs, there is a paradigm shift towards more versatile and general-purpose models that can tackle multiple vision tasks simultaneously. These FMs leverage deep learning techniques, diverse, and large-scale training data to offer a unified framework for various computer vision tasks, enabling researchers to explore new frontiers in visual understanding and analysis (Bommasani et al., 2021). For example, Florence (Yuan et al., 2021), a newly introduced computer vision FM, offered an extensive range of capabilities for visual understanding. By leveraging universal visual-language representations from vast image-text datasets, Florence seamlessly adapted to various computer vision tasks including classification, retrieval, object detection, visual question answering (VQA), image captioning, video retrieval, and action recognition. Notably, Florence achieved outstanding performance in transfer learning scenarios, excelling in fine-tuning, linear probing, few-shot transfer, and zero-shot transfer for novel images and objects. Recently, the Segment Anything Model (SAM, (Kirillov et al., 2023)) was introduced by Meta AI for image segmentation. With its promptability and efficient architecture, SAM achieved impressive zero-shot performance on numerous segmentation tasks (Zhang et al., 2023d, Ma and Wang, 2023, Yang et al., 2023c, Tang et al., 2023). To promote research in FMs for computer vision, the SA-1B dataset, the largest of its kind, providing over 1 billion masks from 11 million images, was open-sourced [^1] SAM project https://segment-anything.com/. Inspired by SAM, the Segment Any Anomaly + (SAA+, (Cao et al., 2023b)) was proposed for zero-shot anomaly segmentation with hybrid prompt regularization to improve the adaptability of modern FMs. Furthermore, in Tang et al. (2023), the authors showed that SAM shows promising performance in camouflaged object detection (COD) tasks, in which objects are seamlessly integrated into their surroundings. In Yang et al. (2023c), SAM demonstrated superior performance in the poultry industry, specifically in the context of segmenting cage-free hens, outperforming SegFormer (Xie et al., 2021) and SETR (Liu et al., 2022) in both whole and part-based chicken segmentation. The Segmenting Everything Everywhere Model (SEEM, (Zou et al., 2023)) represented a significant advancement in the field of image segmentation. It sets a new benchmark for simultaneously segmenting everything everywhere in an image, including semantic, instance, and panoptic segmentation (Kirillov et al., 2019). Notably, SEEM supports a wide range of prompts, including visual, textual, and referring region prompts, which can be combined in various ways to enable versatile and interactive referring segmentation. This model represents a major breakthrough that pushes the boundaries of what is achievable in comprehensive image segmentation tasks.
[…]
Looking ahead, VFMs are expected to continue evolving and pushing the boundaries of what is possible in computer vision. As research progresses, it can be anticipated that such models will be further advanced with improved capabilities in terms of image understanding, scene understanding, and content generation. Furthermore, the fusion of vision models with other modalities, such as language and audio, is likely to lead to the emergence of more powerful and versatile multimodal models (Section 2.2.3).
2.2.3 Multimodal foundation models (MFMs)
Multimodal foundation models (MFMs) represent a new class of artificial intelligence models that leverage and integrate multiple types of data inputs, like text, video, audio, and images, which is a distinct departure from their non-multimodal counterparts that are constrained to processing a singular data type. By capitalizing on the richness and diversity of these data types, MFMs can attain a more holistic understanding of the context. This, in turn, empowers them to execute more nuanced and precise tasks, such as translating text into visual representations, identifying objects in images based on textual prompts, and creating image captions. As a result, these models can fuel a vast array of real-world applications (Bommasani et al., 2021, Moor et al., 2023, Mai et al., 2023).
One notable example is OpenAI’s CLIP (Contrastive Language-Image Pretraining, (Radford et al., 2021)), which has garnered significant attention for its ability to learn joint representations of images and text. It utilizes a strategy known as contrastive learning (Zhao et al., 2021, Jaiswal et al., 2020), which entails learning to associate accurate text-image pairs while dissociating inaccurate ones. Consequently, CLIP excels at tasks such as zero-shot learning, wherein it can classify new images based on textual descriptions and vice versa. Its multimodal nature renders it highly advantageous in diverse domains, such as computer vision, natural language processing, and digital content creation. Building on this foundation, OpenCLIP (Ilharco et al., 2021), a model trained on up to 2B image-text pairs, was proposed to investigate scaling laws (Kaplan et al., 2020) for CLIP (Radford et al., 2021) model with public datasets. OpenCLIP identified power scaling laws through various downstream tasks, such as zero-shot retrieval, classification, linear probing, and end-to-end fine-tuning, concluding that the training distribution plays a key role in scaling laws.
However, a majority of foundation models are typically engineered for either understanding-based or generation-based tasks. To address this issue, BLIP (Li et al., 2022), a new Vision-Language Pre-training (VLP) framework, was introduced for both vision-language understanding and generation tasks. To effectively utilize the noisy web data by bootstrapping the captions, a captioner was employed to generate synthetic captions, subsequently, a filter was adopted to remove the noisy ones. BLIP achieved state-of-the-art results across various vision-language tasks, such as image-text retrieval, image captioning, and VQA. Similarly, OpenFlamingo (Alayrac et al., 2022) was trained on large-scale multimodal web corpora with arbitrarily interleaved text and images for a variety of image and video tasks via few-shot learning, such as image captioning and multiple-choice VQA (Alayrac et al., 2022). KOSMOS-1 (Huang et al., 2023), a Multimodal Large Language Model (MLLM), was subsequently introduced to process multimodal inputs, adhere to instructions, and perform in-context learning for language tasks and multimodal tasks, such as multimodal dialogue, image captioning, VQA, and image recognition with descriptions. Impressively, KOSMOS-1 demonstrated exemplary performance across various settings, such as zero-shot, few-shot, and multimodal chain-of-thought prompting, on various tasks without necessitating any gradient updates or fine-tuning. Built on previous versions, GPT-4 (OpenAI, 2023), a large-scale and multimodal model was introduced by OpenAI to accept both image and text inputs and produce text outputs.
The future of multimodal foundation models is set to be exciting, with further advancements anticipated in several areas. They will likely include more sophisticated integration of various data modalities and enhanced transfer learning for broader applicability across different domains (Bommasani et al., 2021). Furthermore, efforts are expected to be directed towards minimizing data bias from different domains, improving fairness, and developing energy-efficient AI models (Bommasani et al., 2021).
2.2.4 Reinforcement learning foundation models (RLFMs)
Over the past decade, the field of decision making has undergone a significant revolution due to the emergence of reinforcement learning (RL) algorithms integrated with advanced DL/ML algorithms. These models, e.g., deep reinforcement learning (DRL) algorithms like DQN (Mnih et al., 2015) and PPO (Schulman et al., 2017), have enabled RL agents to learn complex behaviors and achieve remarkable performance in various domains, such as game playing (Mnih et al., 2013, 2015, Vinyals et al., 2017), robot control (Kober et al., 2013, Lee et al., 2020), and agricultural applications (Bu and Wang, 2019, Binas et al., 2019, Gandhi, 2022, Gautron et al., 2022). By interacting with the environment, these models learn optimal decision-making through trial-and-error. However, due to their reliance on learning tasks from scratch without comprehensive knowledge of vision, language, or other datasets, these methods often face challenges in terms of generalization and sample efficiency (Yang et al., 2023b).
There are increasing attempts to combing FMs and sequential decision making to tackle complex real-world problems with better generalization. Gato (Reed et al., 2022), a multi-modal, multi-task, and multi-embodiment generalist RL policy, was proposed by Google DeepMind with a 1.2B parameter decoder-only transformer architecture, trained on 604 distinct tasks with varying modalities, observations, and action specifications. Gato was able to play Atari, caption images, chat, and stack blocks with a real robot arm, using the same network and weights with little fine-tuning. Subsequently, AdA (Team et al., 2023) proposed by Google DeepMind with 265M parameters that learned new tasks in an extended XLand 3D environment simulation as fast as humans based on a modified transformer architecture to store more information and enable efficient training.
Leveraging world knowledge from FMs, sequential decision making can solve tasks faster and generalize better (Yang et al., 2023b). FMs for decision making are still in their infancy, still facing significant challenges including the discrepancy in data modalities, uncertainties surrounding environment and task structures, and the absence of certain components in existing FMs and decision-making paradigms. Further advancements are necessary to address these challenges and unlock the full potential of FMs in decision making (Yang et al., 2023b, Reed et al., 2022).
3 Agricultural Foundation Models (AFMs)
In this section, we begin by examining the potential for creating foundation models in agriculture, i.e., agricultural foundation models (AFMs). We then present current applications of these models in the realm of smart agriculture, followed by a discussion on emerging areas where these foundation models are most applicable.
3.1 Developing FMs for Agriculture
Figure 1 presents a comprehensive overview of the pipeline for developing (multimodal) foundation models in agriculture. This process principally encompasses stages such as data collection, dataset curation, model training, and downstream fine-tuning.
The development of multimodal foundation models for agricultural applications commences with the imperative phase of data collection and curation (Bommasani et al., 2021). This process necessitates the assembly of heterogeneous data spanning a range of modalities, including visual (e.g., images and videos indicating the status of crops, pests, and disease symptoms), textual (e.g., human instructions and pest identification documents), meteorological (e.g., temperature, precipitation, and wind speed measurements), and possibly auditory data (e.g., distinctive insect or animal sounds). Given the magnitude of the data involved, meticulous annotation and curation are indispensable to discriminate meaningful patterns and correlations. Such data labeling could encompass tasks such as the classification of crop and weed species, the annotation of disease symptoms present on plant foliage, or the categorization of weather conditions.
Subsequent to data curation is the model training phase. Capitalizing on contemporary advancements in machine learning and artificial intelligence, such as transformer architectures (Vaswani et al., 2017) and convolutional neural networks (Gu et al., 2018), the model is trained to process and interpret multimodal inputs. These artificial intelligence models are designed to associate visual depictions with their corresponding textual labels, comprehend textual descriptions, and generate an inclusive understanding of the agricultural ecosystem. Notably, the training of these models may be computationally demanding, which necessities the use of advanced hardware and optimization algorithms to enhance performance, which will be discussed in Section  4.2.
The culmination of this process is the implementation of these trained models in downstream agricultural applications, such as plant health monitoring (Xu et al., 2022a), crop and weed management (Hu et al., 2021, Dang et al., 2023, Rai et al., 2023), fruit picking (Chu et al., 2021), and precision livestock farming (Yang et al., 2023c). For instance, in relation to plant health, the models can be utilized to identify early indicators of disease or pest infestations, thus facilitating prompt interventions to reduce crop loss. For weed management, the models can assist in distinguishing weeds from crops, enabling precision farming equipment to selectively eradicate unwanted flora. Furthermore, in the domain of precision livestock farming, these models can be harnessed to monitor livestock health and behavior, optimize feeding schedules, and predict potential diseases. In the context of downstream tasks, transfer learning or fine-tuning (Weiss et al., 2016, Zhuang et al., 2020) of the foundation models becomes paramount to adapt the generalized learning to specific agricultural applications, thereby enhancing model performance and efficiency. Through these varied applications, multimodal foundation models hold the potential to significantly transform the agricultural industry, thereby enhancing productivity, promoting sustainability, and ultimately boosting profitability.
However, constructing an agricultural foundation model from scratch presents significant challenges. Firstly, initial data collection requires a large and diverse dataset, which requires laborious efforts and potentially high costs (Lu and Young, 2020). The subsequent data labeling and curation require meticulousness and expertise. Additionally, the process of training models demands substantial computational resources, thus presenting impediments for individuals or entities with constrained resources. Lastly, fine-tuning of the model for specific agricultural applications can be challenging, requiring a balance between generalizability and specificity. A detailed discussion of these challenges will be presented in Section  4.
3.2 Agricultural applications
In this subsection, we explore the existing applications of integrating FMs discussed in Section  2 into the field of smart agriculture.
3.2.1 LFMs in agriculture
Language Foundation Models (LFMs), also referred to as Large-Scale Language Modeling (LLMs) and detailed in Section  2.2.1, have transformed the realms of design, development, and deployment by enabling human-like interactions. Researchers are now increasingly integrating these LFMs into various stages of design and development for agricultural applications, as highlighted in (Stella et al., 2023, Lu et al., 2023). Specifically, in Stella et al. (2023), the authors discussed the process of integrating Large Language Models (LLMs) into the design phase of robotic systems, as depicted in Figure 2 from Stella et al. (2023). They specifically illustrate the procedure to design a robotic gripper optimized for tomato picking. During the initial ideation phase, researchers consult LLMs, such as ChatGPT (OpenAI, 2022), to understand potential challenges and opportunities within the field. Based on this information, they select the most promising and intriguing pathways, subsequently narrowing down the design possibilities through further dialogue with the LLM. This collaborative interaction covers a wide spectrum of knowledge domains and various levels of abstraction, encompassing everything from conceptual thinking to technical implementation. Throughout this process, the human collaborator leverages the AI’s knowledge to access information beyond their personal expertise. During the subsequent phase of the design process, which is more technically focused, these general directions need to be actualized into a tangible, fully functional robot. As of now, LLMs are not capable of producing complete CAD models, assessing code, or autonomously fabricating a robot. However, recent progress in AI research has indicated that these algorithms can provide substantial assistance in executing software (Chen et al., 2021), mathematical reasoning (Wolfram, 2023), and even in the generation of shapes (Ramesh et al., 2022). Although we foresee that future AI methodologies will be capable of handling these tasks, at the moment, the technical implementation remains a team effort between AI models and humans. Finally, the human fine-tuned the code suggested by the LLM, finalized the CAD design, and oversaw the actual fabrication of the robot. Figure 2 (b) from Stella et al. (2023) showed the primary outputs generated by the LLM, as well as the real-world application of the AI-designed robotic gripper used for crop harvesting.
The authors in Lu et al. (2023) have deliberated on the prospective use cases of Large Language Models (LLMs) in organizing unstructured metadata, converting metadata from one format to another, and identifying potential errors during the data collection phase. They also envision the future generation of LLMs as exceptionally powerful tools for data visualization (Bubeck et al., 2023), offering invaluable assistance to researchers in extracting meaningful insights from vast volumes of phenotypic data.
3.2.2 VFMs in agriculture
Figure 3: Side-veiw (a,b) and top-view (c,d) views of chicken segmentation using SAM model in a research cage-free house proposed in Yang et al. (2023c).
Vision Foundation Models (VFMs, as detailed in Section 2.2.2), encompassing algorithms for core computer vision tasks like classification, detection, segmentation, and generative modeling, have revolutionized traditional computer vision methodologies. Trained on extensive image datasets, these VFMs can be efficiently adapted to novel domains, such as agriculture, with minimal fine-tuning and a limited set of labeled images. Recently, the agricultural community has shown a growing interest in VFMs, as seen in recent studies (Yang et al., 2023c, Williams et al., 2023).
In Yang et al. (2023c), the Segment Anything Model (SAM, Section  2.2.2) was employed for chicken segmentation tasks in a zero-shot manner using both part-based segmentation and the use of infrared thermal images (Figure 3). Experimental results showed that SAM outperformed the other VFMs algorithms, such as SegFormer and SETR, in both whole and partial chicken segmentation. For example, SAM achieved a mean Intersection over Union (mIoU) of 94.80% in the chicken segmentation tasks, while SegFormer and SETR only achieved mIoUs of 43.22% and 42.90%, which greatly shows the effectiveness of SAM on the poultry segmentation tasks. In addition, a novel single-bird tracking algorithm, integrating SAM, YOLOX (Ge et al., 2021) and ByteTracker (Zhang et al., 2022b), was proposed for tracking the activities for individual birds across video frames by extracting the location information (i.e., bounding boxes provided by SAM) in each image and then tracking the birds over time with YOLOX and ByteTracker. Despite its promising performance, the tracking algorithm lacked end-to-end capability. The object tracking FMs (Yang et al., 2023a) showcases impressive performance in video object tracking and segmentation, requiring very little human participation, thereby holding the potential to serve as an end-to-end method.
In Williams et al. (2023), an automatic leaf segmentation pipeline, named “Leaf Only SAM”, was introduced for zero-shot segmentation of potato leaves. When compared with a fine-tuned Mask R-CNN (He et al., 2017) model on the annotated potato leaf dataset, Leaf Only SAM demonstrated an average recall and precision of 63.2% and 60.3%, respectively. These results are evidently lower than the Mask R-CNN model’s average recall and precision of 78.7% and 74.7%, respectively. However, it’s important to note that Leaf Only SAM does not necessitate any fine-tuning or an annotated dataset, thereby showing notable zero-shot generalization capabilities. These findings suggest that the direct application of FMs in a zero-shot manner may not always result in high performance due to potential distribution shifts (Chawla et al., 2021). This will be further discussed in Section  4.3.
3.2.3 MFMs in agriculture
Multimodal foundation models (MFMs, Section  2.2.3) introduce a novel class of artificial intelligence models. These leverage and incorporate various types of data inputs including text, video, audio, and images. The agricultural community has begun to delve into multimodal learning within agricultural applications (Bender et al., 2020, Garg et al., 2021, Parr et al., 2021, Cao et al., 2023a, Dhakshayani and Surendiran, 2023) to enhance performance through the utilization of multimodal data. For instance, Bender et al. (2020) open-sourced a multimodal dataset designed for agricultural robotics, which was gathered from cauliflower and broccoli fields, to facilitate robotics and machine learning research activity in agriculture. The dataset includes a variety of data such as stereo color, thermal and hyperspectral imagery, and crucial environmental details like weather and soil conditions. In Cao et al. (2023a), a multi-modal language model based on image-text-label information was proposed for cucumber disease recognition. They effectively merged label information across multiple domains by employing image-text multimodal contrastive learning and image self-supervised contrastive learning. This methodology helped measure the distance of samples within the common image-text-label space. The experimental results revealed that this novel approach attained a recognition accuracy rate of 94.84% on a modestly sized multimodal cucumber disease dataset.
MFMs have also been discussed in the agriculture domain. Lu et al. (2023) investigated the potential of GPT-4, a multimodal language and vision model (as referenced in ChatGPT, Section 2.2.1), in achieving Artificial General Intelligence (AGI). In Stella et al. (2023), Large Language Models (LLMs, e.g., ChatGPT) were utilized in the design phase of robotic systems for tomato picking. However, it’s noteworthy that current applications are predominantly based on text-image data and their use is often confined to question-answering tasks. Indeed, there is a conspicuous gap in applications related to agricultural robotics where inputs such as images, text, or voice (e.g., human instructions), and depth information (e.g., from LiDAR or laser sensors) are used. These robots, which are typically designed for tasks like fruit harvesting or crop monitoring (Tao and Zhou, 2017), could greatly benefit from the application of MFMs, which will be further discussed in Section  3.3.5.
3.2.4 RLFMs in agriculture
Deep reinforcement learning (DRL) (Arulkumaran et al., 2017, Henderson et al., 2018) is an emerging field that fuses deep learning with reinforcement learning to tackle decision-making tasks. Its impressive performance has been demonstrated across a broad spectrum of applications, from computer games (Mnih et al., 2013, 2015), to intelligent transportation systems (Chu et al., 2019, Chen et al., 2023), and robot control (Gu et al., 2017, Lee et al., 2020). Recently, the agricultural community has shown heightened interest in DRL (Binas et al., 2019, Gandhi, 2022), leading to a host of innovative applications such as crop management (Gautron et al., 2022, Din et al., 2022), agriculture robots (Yang et al., 2022b), and smart irrigation systems (Zhou, 2020). However, due to their reliance on learning tasks from scratch without comprehensive knowledge of vision, language, or other datasets, these methods often face challenges in terms of generalization and sample efficiency (Yang et al., 2023b).
Although there currently exist no published papers specifically focusing on the application of reinforcement learning foundation models (RLFMs, Section 2.2.4) in agricultural scenarios, the field nonetheless presents vast untapped opportunities for exploration and innovation. For instance, Gato (Reed et al., 2022) and AdA (Team et al., 2023), discussed in Section 2.2.4, exhibit promising potential for handling decision tasks that involve multimodal, multi-task contexts, paving the way for innovative agriculture applications in the future.
3.3 Emerging opportunities
In this subsection, we discuss some emerging areas in the agricultural space where these foundation models are most applicable.
3.3.1 Smart crop management
To achieve sustainable farming, smart crop management (Eli-Chukwu, 2019, Jaramillo et al., 2020) has evolved as a novel research direction. It integrates advanced methodologies from artificial intelligence, IoT, and big data into crop management tasks, including plant growth monitoring (Lakshmi and Gayathri, 2017), disease detection (Xu et al., 2022a), yield monitoring (Khanal et al., 2018, Kayad et al., 2019), and harvesting (Meshram et al., 2021). Yet, developing such algorithms for precision crop management is complex and challenging, particularly when considering factors like diverse crop varieties, environmental dynamics, and limited or hard-to-acquire training datasets. Furthermore, these systems often necessitate diverse data inputs sourced from varying domains. For example, while a deep learning method is developed on satellite imagery to assess crop health, it simultaneously might need soil sensor data to understand underground moisture levels and market data to predict the optimal time for harvest based on pricing trends. Designing such algorithms requires deep research insights, making it challenging to smoothly incorporate knowledge from unrelated fields.
In light of the aforementioned challenges, Foundation Models (FMs) offer a promising solution. These models, trained on datasets spanning multiple domains, can be seamlessly adapted to new areas with minimal fine-tuning. They hold significant potential as tools for smart crop management. For instance, Language Foundation Models (LFMs, Section  2.2.1) empower researchers to tap into knowledge beyond their personal expertise, facilitating software development and design (Stella et al., 2023). Moreover, multi-modal foundation models (MFMs, Section  2.2.3) integrate visual, textual, and sometimes auditory data, providing a comprehensive analytical lens, proving invaluable in deciphering intricate agricultural situations, notably in precision crop management.
3.3.2 Smart plant breeding
Smart plant breeding, a breakthrough approach, combines data from genotype, environment, and their interaction to optimize crop varieties (Xu et al., 2022b, Zhang et al., 2023c). Employing sophisticated “omics” technologies, artificial intelligence, and big data, it constructs effective models to predict plant traits and performance (phenotype). The objective of this “smart breeding” is to promote genetic gains, shorten breeding cycles, and develop plants that are acclimatized to specific and evolving environments (Xu et al., 2022b). However, the accuracy in phenotype prediction is intrinsically tied to the processing of multidimensional or multimodal data, sourced from spatiotemporal omics, encompassing genotype, environment, and their interaction. Such a data-rich environment poses significant challenges to conventional machine learning methodologies (Xu et al., 2022b).
In response to this complexity, multimodal foundation models (MFMs, Section  2.2.3) may emerge as a potential solution. These models, designed to handle and learn from varied data types across multiple domains, could significantly improve the accuracy and efficiency of predicting plant phenotypes, thereby pushing the field of smart plant breeding toward unprecedented possibilities.
3.3.3 Smart livestock farming
Smart livestock farming is an innovative, automated approach that fuses cutting-edge technologies such as the Internet of Things (IoT), deep learning, and technology for intelligent control of livestock production (Islam et al., 2019, Astill et al., 2020, Zhang et al., 2021b, Farooq et al., 2022). This includes the automation of critical processes such as food and water supply, egg collection, livestock activity monitoring, and maintenance of precise environmental conditions. The primary objective of this innovative methodology is to enhance operational efficiency in livestock farms, leading to considerable savings in labor, maintenance, and material costs (Choukidar and Dawande, 2017). Smart livestock farming, while innovative and promising, also presents challenges. The development and implementation of such an intelligent system require a multidisciplinary understanding that spans the realms of advanced technologies, animal science, environmental science, and data analytics, among others (Islam et al., 2019). This integrative approach involves handling and processing voluminous and varied data across multiple domains. Overcoming these challenges would require strategic collaborations, expertise across domains, and harnessing cutting-edge tools in AI and data analytics.
Given these complex challenges, FMs emerge as a potential solution to facilitate the realization of smart livestock farming (Yang et al., 2023c). These comprehensive models, built on large-scale multimodal data and incorporated knowledge from various domains, can effectively manage the complications of multi-dimensional information and bridge the gaps between different disciplines (Bommasani et al., 2021). FMs, with their ability to learn and generalize from large datasets, could provide valuable insights, enhance decision-making, and enable more precise management, thereby contributing significantly to the advancement of intelligent and efficient livestock farming systems.
3.3.4 Precision aquaculture farming
Precision aquaculture farming (O’Donncha and Grant, 2019, Føre et al., 2018, Antonucci and Costa, 2020), also known as smart aquaculture or precision fish farming, is a modern approach to fish and shellfish farming that integrates advanced technologies like IoT and data-driven techniques to optimize production efficiency, environmental sustainability, and animal welfare. It is often characterized by hundreds of interconnected sensors that store and serve data to gather essential information such as temperature, oxygen, weather, chlorophyll, fish tags, imagery data, and 3D sensor data, which can be utilized for monitoring fish variables (species, health, etc.), behavioral analysis, feeding decisions, and water quality prediction (Yang et al., 2021). Recently, deep learning (DL) has emerged as a promising solution to address the problems in precision aquaculture farming. For example, in Måløy et al. (2019), a dual-stream recurrent network was adopted to classify fish feeding status for salmon, yielding 80% feeding classification accuracy. However, these DL approaches rely heavily on large-scale and labeled datasets for training (Yang et al., 2021). Underwater conditions can be adverse, with poor illumination and low visibility in turbid water, as well as cluttered backgrounds, making it difficult to acquire high-quality data. Besides, it is very challenging to integrate, handle, and process numerous and varied data across multiple domains.
Fortunately, FMs have become an encouraging solution to facilitate the realization of precision aquaculture farming, representing the unique capability to handle diverse data from multiple domains through few-short or zero-short learning. For example, generation-based FMs can be harnessed for data creation, while RLFMs can be utilized for decision-making in fish farming. We firmly believe that the application of FMs could significantly enhance the accuracy and efficiency of associated applications in precision aquaculture farming.
3.3.5 Agriculture robots
Agricultural robots, often referred to as “AgriRobots”, are at the forefront of transforming modern agriculture (Bechar and Vigneault, 2016). AgriRobots can automate various farming tasks such as planting, weeding, harvesting, and crop health monitoring (Fountas et al., 2020). For instance, AI-guided robots can identify and remove weeds with precision, significantly reducing the use for herbicides. Similarly, drones equipped with advanced imaging sensors can monitor fields for signs of disease or pests, enabling early intervention and reducing yield losses. Similar to the contexts of smart breeding and smart livestock farming, the development and implementation of AgriRobots require multidisciplinary knowledge including robotics, AI, plant science, environmental science, and data analytics. This mixed approach involves handling and processing a significant amount of data from various domains, making it a complex endeavor (Bechar and Vigneault, 2016, Fountas et al., 2020).
However, these challenges can be guided effectively by the potential application of FMs. These models, trained on large-scale multimodal data, can provide the robust predictive capabilities, which are necessary for advanced farming operations. For instance, Reinforcement Learning Foundation Models (RLFM, Section  2.2.4) could enable precision control of AgriRobots (Reed et al., 2022, Team et al., 2023), while Language Foundation Models (LFM, Section  2.2.1) could provide valuable insights during the design process of AgriRobots (Stella et al., 2023). By incorporating insights from diverse domains, FMs can offer an integrated solution, driving the development of more effective and efficient agricultural robots, and thus, revolutionizing the future of agriculture.
4 Challenges and Outlook
Although FMs have demonstrated impressive performance in various agriculture applications, there remain open challenges with training, deployment, and distribution shifts, as identified and discussed below.
4.1 Data collection
The process of data collection for training FMs in agricultural applications presents unique challenges due to the inherent variability and diversity in agriculture, influenced by factors such as crop types, growth stages, soil conditions, weather patterns, farming practices, etc. These factors result in a complex and highly diverse dataset, making its collection and standardization a challenging task (Lu and Young, 2020). Firstly, it is critical to ensure the collected data is accurate, comprehensive, and adequately captures the wide variability found in different farming environments or diverse range of growth phases. However, acquiring high-quality data can be time-consuming, labor-intensive, and expensive, especially considering the need for ground truth labeling in supervised learning scenarios (Li et al., 2023a) and the variety growth stages of plants or livestocks. To tackle these challenges, generative models (Lu et al., 2022) that create high-fidelity images and label-efficient learning methods to save on tedious labeling costs (Li et al., 2023a, Zhang et al., 2023e) show promises. Secondly, data privacy and ownership pose challenges as farms are typically privately owned. Farmers may be reluctant to share data due to concerns over privacy or potential commercial exploitation. Novel frameworks, such as federated learning (Zhang et al., 2021a), can address these concerns, paving the way for broader access to high-quality, diverse agricultural datasets. Temporal dynamics represent another significant hurdle. Agricultural processes inherently change over time, influenced by daily fluctuations, seasonal variations, or annual changes. This requires the collection of time-series data (e.g., data across years), introducing another layer of complexity. To handle temporal dynamics, techniques such as time-series analysis (Ostrom, 1990), dynamic models based on Long Short-Term Memory (LSTM) (Graves and Graves, 2012), and data augmentation (Lu et al., 2022) can be employed to better account for the inherent temporal dynamics in agricultural data.
In conclusion, while data collection for training FMs in agriculture is a challenging endeavor, addressing these issues is crucial for developing robust and effective models capable of advancing the state of agricultural technology and practice.
4.2 Training efficiency
Training FMs for agriculture applications presents significant challenges, particularly with respect to long training times and substantial costs, e.g., requiring thousands of GPU hours and millions of training data. However, innovative solutions are being devised to overcome these challenges. One such approach is the integration of an existing LLM with a relatively lightweight visual prompt generator (VPG) (Li et al., 2023b). Though this approach can reduce the computational costs, the tuning of the VPG component within the VL-LLM still represents a significant computational demand (Zhang et al., 2023a). A promising solution discussed in the literature involves the transfer of an existing VPG from any existing FMs to the target FMs. This allows for the leveraging of pre-existing model structures and parameters, thereby bypassing the need for extensive training from scratch. The key to optimizing this process lies in maximizing transfer efficiency across different LLM sizes and types. A significant development in this regard is the VPGTrans approach (Zhang et al., 2023a), a two-stage transfer framework designed to speed up the transfer learning process, which demonstrated that their method can facilitate VPG transfer with a speed-up of more than 10×10$\times$10 × and uses only a fraction of the training data compared to building a VPG from scratch.
In summary, while training FMs for agricultural applications indeed presents substantial challenges due to long training times and high costs, advancements like the VPG transfer technique offer promising solutions. These techniques pave the way for more efficient and cost-effective model training processes, thereby making the use of FMs more accessible and practical in the field of agriculture.
4.3 Distribution shift
The issue of distribution shift presents a significant challenge when applying FMs to agricultural applications. Distribution shift occurs when the data that the model encounters during deployment differs significantly from the data used during its training phase (Wiles et al., 2022). In the context of agriculture, the environmental conditions under which data is collected can vary greatly across different regions and seasons. These variations can include differences in crop types, soil conditions, weather patterns, and farming practices, all of which can cause considerable shifts in data distribution (Wiles et al., 2022). For example, Williams et al. (2023) demonstrated that the direct application of FMs in a zero-shot manner for leaf segmentation tasks resulted in unsatisfactory performance, likely due to potential distribution shifts (Chawla et al., 2021).
To combat the challenges posed by distribution shifts when training and deploying FMs, researchers have been continuously investigating these issues (Chawla et al., 2021, Cvejoski et al., 2022, Bommasani et al., 2021). For instance, Cvejoski et al. (2022) demonstrated that Large Language Models (LLMs) may experience performance drops of approximately 88% at best when predicting the popularity of future posts from subreddits whose topic distribution evolves over time. To mitigate these distribution shift issues, they introduced neural variational dynamic topic models and attention mechanisms. These techniques infer temporal language model representations for regression tasks, resulting in only about 40% performance drops in the worst-case scenarios and a mere 2% in the best cases. These techniques have been shown to significantly reduce performance drops caused by distribution shifts, thus demonstrating promising prospects for the application of FMs in the diverse and dynamic field of agriculture.
4.4 Other challenges in real-world deployment
Deploying FMs for agricultural applications presents a unique set of challenges, particularly regarding inference speed and model size. Firstly, the inherent complexity and size of these models can pose substantial computational demands. Large-scale FMs often require significant processing power and memory for both training and inference, which may not always be readily available in on-field agricultural settings (Bommasani et al., 2021). This can limit the deployment of these models in real-time applications, such as monitoring crops or automating irrigation systems. Secondly, inference speed is crucial for many agricultural tasks that require immediate action based on the model’s outputs (Bommasani et al., 2021). For instance, in precision agriculture, time-sensitive tasks like pest detection or predicting weather impacts on crop health demand quick and accurate responses. However, due to the size and complexity of foundation models, achieving high-speed inference can be a challenge.
To overcome these hurdles, future research and development efforts need to focus on model optimization techniques (Zhong et al., 2023), such as model compression (e.g., pruning and quantization (Cheng et al., 2017, Choudhary et al., 2020)) and efficient network design (Wan, 2022), which could potentially reduce the model size without compromising performance. Actually, there are already some attempts for deploying FMs. For instance, a fully quantized vision transformer, FQ-ViT (Lin et al., 2022), was proposed to simplify inference by using only 4-bit quantization and the BitShift operator with only nearly lossless accuracy degradation ($\sim 1\%$). SqueezeLLM (Kim et al., 2023) showed that 3-bit quantization significantly reduces the perplexity gap from the FP16 baseline by up to 2.1 times and achieved up to 2.3 times speedup compared to the baseline models.
Additionally, the utilization of edge computing solutions (Chen and Ran, 2019) could expedite the inference process by enabling data processing closer to its source, consequently enhancing the speed and efficiency of these models in agricultural applications. Recently, MLC LLM (Girshick et al., 2023) offers a universal solution, allowing any language model to be deployed natively across a diverse range of hardware backends and native applications. This solution also provides a productive framework for individuals to further optimize model performance for their specific use cases. On the other hand, MobileSAM (Zhang et al., 2023b) made SAM (Kirillov et al., 2023) mobile-friendly by replacing the heavyweight image encoder with a lightweight one. This lightweight SAM is more than 60 times smaller yet performs on par with the original SAM (Kirillov et al., 2023).
5 Summary
In conclusion, this study has investigated the potential of foundation models in the field of smart agriculture. These models hold the promise of transforming the agricultural sector by reducing reliance on extensive labeled datasets and enhancing generalization and effectiveness. By categorizing and reviewing various types of foundation models (Section 2), we have revealed their applicability in smart crop management, smart plant breeding, smart livestock farming , precision aquacuture farming , and agricultural robotics (Section 3). However, their utilization in smart agriculture is still in its early stages, and challenges and risks such as model training, validation, and deployment need careful consideration (Section 4). Further exploration and utilization of foundation models have the potential to enhance productivity, sustainability, and decision-making in agriculture. This study aims to inspire and guide researchers to unlock the untapped potential of foundation models in smart agriculture. We hope that this manuscript can serve as a stepping stone toward developing smart agriculture with next-level intelligence and more sophisticated capabilities.
Authorship Contribution
Jiajia Li: Conceptualization, Investigation, Software, Writing – original draft; Mingle Xu: Conceptualization, Investigation, Writing – original draft; Lirong Xiang: Investi-
gation, Writing - review & editing; Dong Chen: Conceptualization, Investigation, Software, Writing – original draft; Weichao Zhuang: Writing - review & editing; Xunyuan Yin: Writing - review & editing; Zhaojian Li: Supervision, Writing - review & editing.